OpenStudy (anonymous):

find the exact value of the given trigonometric function csc13π/4

OpenStudy (anonymous):

csc = 1/sin sin(3pi/4 = sqrt(2)/2 csc 3pi/4 = 2/sqrt(2) = sqrt(2)

OpenStudy (anonymous):

oh wait, it says 13 my bad...

OpenStudy (anonymous):

csc = 1/sin sin(13pi/4) = sin (pi/4) = sqrt(2)/2 csc 13pi/4 = 2/sqrt(2) = sqrt(2) regardless its the same thing